# LongVILA: Scaling Long-Context Visual Language Models for Long Videos

### v5

@@ Abstract
Long-context capability is critical for multi-modal foundation models, especially for long video understanding. We introduce LongVILA, a full-stack solution for long-context visual-language models by co-designing the algorithm and system. For model training, we upgrade existing VLMs to support long video understanding by incorporating two additional stages, i.e., long context extension and long video supervised fine-tuning. However, training on long video is computationally and memory intensive. We introduce the long-context Multi-Modal Sequence Parallelism (MM-SP) system that efficiently parallelizes long video training and inference, enabling 2M context length training on 256 GPUs without any gradient checkpointing. LongVILA efficiently extends the number of video frames of VILA from 8 to 2048, improving the long video captioning score from 2.00 to 3.26 (out of 5), achieving 99.8% accuracy in 6,000-frame (more than 1 million tokens) video needle-in-a-haystack. LongVILA-7B demonstrates strong accuracy on the VideoMME benchmark, i.e., 61.8% with subtitle. Besides, MM-SP is 2.1×2.1$\times$2.1 × - 5.7×5.7$\times$5.7 × faster than ring style sequence parallelism and $\times1$ 1.4×1.4$\times$1.4 × faster than Megatron with a hybrid context and tensor parallelism. Moreover, it seamlessly integrates with Hugging Face Transformers.
@@ 1 Introduction
While some works have enabled long-context Vision-Language Models (VLMs) (Lin et al., 2023b; Weng et al., 2024), they employ simplified approaches rather than offering a comprehensive solution. For instance, LongVA (Zhang et al., 2024a) relies on long-context LLMs and trains models on short-context data. LongVLM (Weng et al., 2024) utilizes token compression to circumvent context extension. These approaches sidestep more challenging issues, such as the development of a robust long-context multi-modal training framework and corresponding dataset design.
@@
Figure 1: The LongVILA training pipeline. In Stages 1 through 3, the process starts with alignment, pre-training, and supervised fine-tuning. In Stage 4, the model undergoes mid-training context extension. Finally, in Stage 5, the model is fine-tuned for long video understanding with Multi-Modal Sequence Parallelism (MM-SP). While this example uses VILA, the process is adaptable to other VLMs. We train LongVILA models to support 1024 video frames in Stage 5 in experiments.
Figure 2: Scaling the context length can enhance multi-modality accuracy, improving performance on both Video-QA tasks with VideoMME and video captioning with LongVILA-Caption.
@@
LongVILA-7B presents strong performance on VideoMME (Fu et al., 2024a), 61.8% with subtitle. The LongVILA model, trained on 2048 frames, achieves 99.8% accuracy in the needle-in-a-haystack experiments with 6,000 frames, with a context length of more than 1 million tokens. In ablations, by increasing the number of video frames using LongVILA, the performance on VideoMME and long video captioning tasks consistently improves (Figure 2). Our MM-SP system can efficiently scale the context length up to 2 million tokens without gradient checkpointing, achieving 2.1$\times$ to 5.7$\times$ speedup compared to ring style sequence parallelism, and 1.1$\times$ to 1.4$\times$ compared to Megatron with a hybrid context parallelism and tensor parallelism.
2 Background and Related work
@@ 3.1 Stage1&2&3: Alignment, Pre-training, and Short Supervised Fine-tuning
We first use open-sourced image and video caption datasets to train the multi-modal projector in stage (1) to conduct the multi-modal alignment. Note that, following (Lin et al., 2023b), both vision encoder and LLM decoder are frozen at this stage. After that, we conduct large-scale pre-training to learn general multi-modal capacity at scale. To improve the quality of large open-sourced datasets, we follow VILA2 (Fang et al., 2024) to relabel COYO-25M (Lin et al., 2023b; Byeon et al., 2022) with VILA-1.5-40B (Lin et al., 2023b). The supervised fine-tuning process incorporates mixed data types, including both images and videos. For short video comprehension, we utilize open-source video instruction-following datasets, e.g., YouCook2 Zhou et al. (2018) and ShareGPTVideo Zhang et al. (2024b). In experiments, our model is based on Qwen2-7B (qwe, 2024) for benchmark comparisons, e.g., VideoMME, and Llama3-8B (AI@Meta, 2024) for ablation study.
Figure 5: The pipeline for generating instruction-following data from long videos. The process begins by segmenting the long video into short clips, each approximately 10 seconds in length. These clips are individually annotated with captions using the VILA-1.5 model. Subsequently, a Large Language Model (LLM) is employed to generate question-and-answer pairs based on the captions of these clips. The types of questions generated include summarization and other inquiries pertinent to the content of long videos.
@@ 4.1 Limitations of Existing Systems
In this section, we address the inefficiencies of current sequence-parallel systems when applied to multi-modal LLMs.
@@ 4.3 Multi-Modal Sequence Parallelism Inference Mode
5 Evaluation
We evaluate our full-stack solution from both system and modeling perspectives. First, we present our training and inference results to demonstrate the efficiency and scalability of our system, which enables long-context training and inference. We then assess the performance of our long-context model on captioning and instruction-following tasks.
@@ 5.1.1 Training system
Table 2: Iteration time on the ShareGPT4Video dataset (Chen et al., 2024a) with and without our two-stage sharding strategy, measured under various settings as detailed in §$\S$ 5.1.3 (seconds).
2 GPUs 4 GPUs 8 GPUs
one-stage 0.78 0.89 1.20
two-stage (our MM-SP) 0.77 0.86 1.12
@@ 5.1.3 Effect of two-stage sharding
Table 3: Performance comparison on VideoMME (Fu et al., 2024a) benchmark.
@@
LongVILA 7B 128 57.5 69.3 56.1 47.0 60.6 70.4 59.2 52.1
256 57.4 69.0 56.7 46.4 61.8 70.8 60.6 54.0
Table 4: Ablation on Stage 3 (S-3) and Stage 4 (S-4) for LongVILA on VideoMME.
S-3 S-4 Average Short Medium Long
28.5% 29.2% 28.8% 27.4%
44.9% 55.1% 43.0% 36.5%
49.2% 60.2% 48.2% 38.8%
@@
5.2 General Video Understanding
Table 3 presents the performance of LongVILA, comparing to state-of-the-art models (Lin et al., 2023a; Zhang et al., 2024c; Chen et al., 2024a; Li et al., 2023c; Cheng et al., 2024; Jin et al., 2023; Zhang et al., ; Share, 2024; Chen et al., 2024c; Fu et al., 2024b; Fei et al., 2024; Liu et al., 2024c) on the VideoMME benchmark, comparing their effectiveness across short, medium, and long video lengths, as well as overall performance. LongVILA, utilizing 128 frames, achieves an overall score of 57.5 / 60.6 without / with subtitle, which are competitive results. We also do the ablation on the effects of Stage 3 and Stage 4 in Table 5, using Llama3 based VILA models.
5.3 Needle-in-a-Haystack
In Figure 3, we present the results of the ”Needle in a Haystack” experiment for long videos. Following the methodology established in the existing literature (Zhang et al., ), we prepared a long video and sampled a fixed number of frames. We inserted specifically designed images at various depths and tasked the model with answering corresponding questions. The 32-frame baseline model (left) was unable to accurately retrieve the correct images beyond 32 frames. In contrast, the LongVILA model (right) demonstrated enhanced performance across a range of frame counts and depths.
5.4 LongVILA-Caption
6 Conclusion
We introduce LongVILA, a comprehensive full-stack solution for long-context visual language models, encompassing model training pipeline and distributed system. Based on our curated long video datasets and five-stage training pipeline, our LongVILA model extends the feasible frame count from 8 to 2048, precisely capturing fine-grained information from 2-hour needle-in-a-haystack videos and delivering promising results in both video QA and captioning tasks. Our system efficiently scales context length up to 2 million tokens, achieving speedups of 2.1$\times$ to 5.7$\times$ compared to ring sequence parallelism and 1.1$\times$ to 1.4$\times$ compared to a hybrid Megatron context and tensor parallelism.
Appendix A Appendix
@@
Table 8: Comparison with state-of-the-art methods Li et al. (2023b); Dai et al. (2023); Bai et al. (2023); Liu et al. (2023b); Lin et al. (2023b); Liu et al. (2024b); Tong et al. (2024); Li et al. (2024) on 10 image based VLM benchmarks. S3 refers to the stage 3 model in LongVILA training pipeline.

### v6

@@ Abstract
Long-context capability is critical for multi-modal foundation models, especially for long video understanding. We introduce LongVILA, a full-stack solution for long-context visual-language models by co-designing the algorithm and system. For model training, we upgrade existing VLMs to support long video understanding by incorporating two additional stages, i.e., long context extension and long video supervised fine-tuning. However, training on long video is computationally and memory intensive. We introduce the long-context Multi-Modal Sequence Parallelism (MM-SP) system that efficiently parallelizes long video training and inference, enabling 2M context length training on 256 GPUs without any gradient checkpointing. LongVILA efficiently extends the number of video frames of VILA from 8 to 2048, achieving 99.8% accuracy in 6,000-frame (more than 1 million tokens) video needle-in-a-haystack. LongVILA-7B demonstrates strong accuracy on 9 popular video benchmarks, e.g., 65.1% VideoMME with subtitle. Besides, MM-SP is 2.1×2.1$\times$2.1 × - 5.7×5.7$\times$5.7 × faster than ring style sequence parallelism and $\times1$ 1.4×1.4$\times$1.4 × faster than Megatron with a hybrid context and tensor parallelism. Moreover, it seamlessly integrates with Hugging Face Transformers. Our code and models are available at github.com/NVlabs/VILA/longvila.
@@ 1 Introduction
While some works have enabled long-context Vision-Language Models (VLMs) (Lin et al., 2023b; Weng et al., 2024), they employ simplified approaches rather than offering a comprehensive solution. For instance, LongVA (Zhang et al., 2024b) relies on long-context LLMs and trains models on short-context data. LongVLM (Weng et al., 2024) utilizes token compression to circumvent context extension. These approaches sidestep more challenging issues, such as the development of a robust long-context multi-modal training framework and corresponding dataset design.
@@
Figure 1: The LongVILA training pipeline. In Stages 1 through 3, the process starts with alignment, pre-training, and supervised fine-tuning. In Stage 4, the model undergoes mid-training context extension. Finally, in Stage 5, the model is fine-tuned for long video understanding with Multi-Modal Sequence Parallelism (MM-SP).
@@
Figure 3: Scaling video frames improves VideoMME accuracy in long category.
Training Stages VideoMME
Average Short Medium Long
1-2-3-4-5 57.5 69.3 56.1 47.0
1-2-4-(3&5) 55.9 67.4 54.1 46.1
4-1-2-3-5 56.0 69.2 54.1 44.5
4-1-2-(3&5) 55.3 67.2 53.6 45.1
Table 1: Ablations on various training stage settings on VideoMME (without subtitle). (3&5) means joint training of the stage 3 and 5.
@@
LongVILA-7B presents strong performance on 9 popular benchmarks, e.g., 65.1% on VideoMME (Fu et al., 2024a) with subtitle. The LongVILA model, trained on 2048 frames, achieves 99.8% accuracy in the needle-in-a-haystack experiments with 6,000 frames, with a context length of more than 1 million tokens. In ablations, by increasing the number of video frames using LongVILA, the performance on VideoMME in long videos consistently improves (Figure 3). Our MM-SP system can efficiently scale the context length up to 2 million tokens without gradient checkpointing, achieving 2.1$\times$ to 5.7$\times$ speedup compared to ring style sequence parallelism, and 1.1$\times$ to 1.4$\times$ compared to Megatron with a hybrid context parallelism and tensor parallelism.
2 Related works
@@ 3.1 Stage1&2&3: Alignment, Pre-training, and Short Supervised Fine-tuning
We first use open-sourced image and video caption datasets to train the multi-modal projector in stage (1) to conduct the multi-modal alignment. Note that, following (Lin et al., 2023b), both vision encoder and LLM decoder are frozen at this stage. After that, we conduct large-scale pre-training to learn general multi-modal capacity at scale. To improve the quality of large open-sourced datasets, we follow VILA2 (Fang et al., 2024) to relabel COYO-25M (Lin et al., 2023b; Byeon et al., 2022) with VILA-1.5-40B (Lin et al., 2023b). The supervised fine-tuning process incorporates mixed data types, including both images and videos. For short video comprehension, we utilize open-source video instruction-following datasets, e.g., YouCook2 Zhou et al. (2018) and ShareGPTVideo Zhang et al. (2024c). In experiments, our model is based on Qwen2-1.5B and Qwen2-7B (qwe, 2024).
@@
Figure 5: The pipeline for generating instruction-following data from long videos. The process begins by segmenting the long video into short clips, each approximately 10 seconds in length. These clips are individually annotated with captions using the VILA-1.5 model. Subsequently, an LLM is employed to generate question-and-answer pairs based on the captions of these clips. Generated questions include summarization and other inquiries pertinent to the content of long videos.
@@ 4.3 Multi-Modal Sequence Parallelism Inference Mode
5 Experimental Results
@@ 5.1.1 Training system
Table 3: Performance comparisons on 9 video benchmarks, including ActivityNet-QA (Yu et al., 2019), EgoSchema (Mangalam et al., 2023), EventBench (Du et al., 2024), LongVideoBench (Wu et al., 2024), PerceptionTest (Patraucean et al., 2023), MVBench (Li et al., 2024b), NExT-QA (Xiao et al., 2021), VNBench (Zhao et al., 2024), and VideoMME (Fu et al., 2024a).
Model
LLM Size
ActNet-QA
EgoSchema
EventBench
LVideoBench
PercepTest
MVBench
NExT-QA
VNBench
VideoMME
test test val val val test mc val w/o sub. w/ sub.
GPT-4V - 57.0 - 32.6 61.3 - 43.5 - - 59.9 63.3
GPT-4o - - - 53.3 66.7 - - - 64.4 71.9 77.2
Gemini-1.5-Pro - 57.5 72.2 43.2 64.0 - - - 66.7 75.0 81.3
Video-LLaVA 7B 45.3 38.4 5.9 37.6 - 43.5 - 12.4 39.9 41.6
Flash-VStream 7B 51.9 - - - - - 61.6 - - -
ShareGPT4Video 8B 50.8 - - 41.8 - 51.2 - - 39.9 43.6
VideoLLaMA2 7B 50.2 51.7 6.9 - 51.4 54.6 - 4.5 47.9 50.3
VideoLLaMA2.1 7B 53.0 53.1 - - 54.9 57.3 - - 54.9 56.4
Kangaroo 8B - 62.7 - 54.8 - 61.1 - - 56.0 57.6
PLLaVA 7B 56.3 - 28.2 39.2 - 46.6 - - - -
LLaVA-OV 7B 56.7 60.1 - 56.4 57.1 56.7 79.4 51.8 58.2 61.5
LongVILA 7B 59.5 67.7 58.0 57.1 58.1 67.1 80.7 63.0 60.1 65.1
@@ 5.1.3 Effect of two-stage sharding
Table 4: Performance comparison on VideoMME (Fu et al., 2024a) benchmark in details.
@@
LongVILA 1.5B 256 53.6 66.2 49.3 45.3 57.5 70.2 54.1 48.2
7B 60.1 69.0 58.3 53.0 65.1 72.9 64.9 57.4
5.2 General Video Understanding
Table 3 presents the performance of LongVILA, comparing to state-of-the-art models (Lin et al., 2023a; Zhang et al., 2024d; Chen et al., 2024a; Li et al., 2023c; Cheng et al., 2024; Jin et al., 2023; Zhang et al., ; Share, 2024; Chen et al., 2024c; Fu et al., 2024b; Fei et al., 2024; Liu et al., 2024c; Xu et al., 2024; Li et al., 2024a; Zhang et al., 2024a) on popular video benchmarks. LongVILA-7B achieves strong performance across all these benchmarks. Table 4 compares their effectiveness across short, medium, and long video lengths, as well as overall performance on VideoMME. LongVILA, utilizing 256 frames, achieves an overall score of 60.1 / 65.1 without / with subtitle, which are competitive results. We include LongVILA-1.5B (starting from Qwen2-1.5B (qwe, 2024)) for evaluation, which is also competitive. We provide a detailed model complexity of LongVILA among various model size, number of frames, context length, latency and FLOPs in Table 10 in the appendix. We also do the ablation on training schedules in Table 1.
5.3 Needle-in-a-Haystack
In Figure 2, we present the results of the Needle in a Haystack experiment for long videos. Following the methodology established in the existing literature (Zhang et al., ), we prepared a long video and sampled a fixed number of frames. We inserted specifically designed images at various depths and tasked the model with answering corresponding questions. The 32-frame baseline model (left) was unable to accurately retrieve the correct images beyond 32 frames. In contrast, the LongVILA model (right) demonstrated 99.8% accuracy across 6,000 frames, which contains more than 1 million tokens. To our best knowledge, this is the first VLM which can handle 1 million context length.
6 Conclusion
We introduce LongVILA, a comprehensive full-stack solution for long-context visual language models, encompassing model training pipeline and distributed system. Based on our curated long video datasets and five-stage training pipeline, our LongVILA model extends the feasible frame count from 8 to 2048, precisely capturing fine-grained information from 2-hour needle-in-a-haystack videos, with more than 1 million tokens Our LongVILA-7B model achieves strong performance across popular video benchmarks, especially on VideoMME, e.g., 65.1% accuracy with subtitle. Our system efficiently scales context length up to 2 million tokens, achieving speedups of 2.1$\times$ to 5.7$\times$ compared to ring sequence parallelism and 1.1$\times$ to 1.4$\times$ compared to a hybrid Megatron context and tensor parallelism.
Appendix A Appendix
A.1 LongVILA-Caption
Table 5: Iteration time (seconds) on the dataset (Chen et al., 2024a) with and without our two-stage sharding strategy.
2 GPUs 4 GPUs 8 GPUs
one-stage 0.78 0.89 1.20
two-stage 0.77 0.86 1.12
@@ 2 GPUs 4 GPUs 8 GPUs
Table 9: Comparison with state-of-the-art methods (Li et al., 2023b; Dai et al., 2023; Bai et al., 2023; Liu et al., 2023b; Lin et al., 2023b; Liu et al., 2024b; Tong et al., 2024; Li et al., 2024c) on 10 image based VLM benchmarks. S3 refers to the stage 3 model in LongVILA training pipeline.
@@
Table 10: Detailed model complexity analysis of LongVILA among various model size, number of frames, parameters, latency (ms), and TFLOPs. We profile LongVILA models into 4 types of components. Image Encoder includes vision tower and mm projector. LLM Linears include k/q/v/o projectors and linears in decoder layers. LLM Attention is the attention computation. LLM Others include other components, like embeddings, output heads, and normalization layers. We use fp16 data type, Flash-Attention2 (Dao, 2024) on one A100 GPU for latency measurement. As the number of frames increases, the FLOPs and latency of LLM Attention grow quadratically, whereas other components increase linearly. LLM Attention represents the predominant computational cost in long video understanding, highlighting the MM-SP system.
Frames Context Metric LongVILA-1.5B LongVILA-7B
Image
Encoder
LLM
Linears
LLM
Attention
LLM
Others
Image
Encoder
LLM
Linears
LLM
Attention
LLM
Others
Params 0.44B 1.31B - 0.23B 0.46B 6.53B - 1.09B
32 6415 Latency 196.0 109.3 27.0 28.9 201.0 426.2 51.1 42.1
TFLOPs 5.13 4.20 1.77 0.75 5.22 20.93 4.13 1.74
64 12719 Latency 375.2 208.1 80.7 50.9 375.8 831.1 173.8 80.8
TFLOPs 10.26 8.33 6.96 1.48 10.45 41.50 16.24 3.46
128 25327 Latency 740.4 403.2 288.1 97.1 755.1 1642.4 644.9 158.3
TFLOPs 20.52 16.60 27.59 2.95 20.89 82.64 64.38 6.89
256 50543 Latency 1456.0 811.2 1087.7 192.5 1476.3 3308.2 2529.6 325.1
TFLOPs 41.05 33.12 109.89 5.88 41.78 164.92 256.40 13.74
512 100975 Latency 2921.0 1660.9 4359.2 389.4 2980.2 6675.1 10149.3 653.4
TFLOPs 82.09 66.17 438.59 11.75 83.57 329.48 1023.37 27.45